OpenStudy (anonymous):

Calculate the change internal energy (ΔE) for a system that is giving off 45.0 kJ of heat and is performing 855 J of work on the surroundings.

OpenStudy (anonymous):

I thought it would be -44.1 but the answer key says -45.9

OpenStudy (anonymous):

try using mcat and the rest of the question

OpenStudy (aaronq):

giving off heat is negative and so is doing work. \(\Delta U=q+w=-45000~J+(-855~J)\)

OpenStudy (photon336):

If work is done on the system it's positive; if it's done to on the surroundings it's negative. if heat leaves the system q <0 if heat enters system q >0 total internal energy delta E = q+ w